A baseball is the thrown at 30 m/s. The baseball has mass of 0.5 kg. The ball was in the hand of the pitcher for 0.3 seconds. How large a force did the pitcher exert on the ball?
1
Expert's answer
2020-11-12T09:33:49-0500

We can find the average force that the pitcher exert on the ball from the impulse-momentum change equation:


"F_{avg}\Delta t=m\Delta v=m(v_f-v_i),"

here, "F_{avg}" is the average force that the pitcher exert on the ball, "\Delta t=0.3\ s" is the time during that the ball was in the hand of the pitcher, "m=0.5\ kg" is the mass of the ball, "\Delta v" is the change in the velocity of the ball, "v_i=30\ \dfrac{m}{s}" is the initial velocity of the ball, "v_f=0\ \dfrac{m}{s}" is the final velocity of the ball.

Then, from this equation we can find the magnitude of the average force that the pitcher exert on the ball:


"F_{avg}=\dfrac{m \lvert (v_f-v_i) \rvert}{\Delta t},""F_{avg}=\dfrac{0.5\ kg\cdot \lvert (0\ \dfrac{m}{s}-30\ \dfrac{m}{s}) \rvert}{0.3\ s}=50\ N."

Answer:

"F_{avg}=50\ N."